OpenStudy (anonymous):

\[\int\limits_{1}^{3} e^{3x} dx\]

hartnn (hartnn):

start with substituting 3x = u du =.... ?

hartnn (hartnn):

@Jlockwo3 can you find du=.... ?

OpenStudy (anonymous):

is it 1/3 du

hartnn (hartnn):

i was asking du in terms of dx u= 3x differentiate, w.r.t x., what u get ?

OpenStudy (anonymous):

guess idk what the question is, im lost

hartnn (hartnn):

but you can differentiate, right ? what you get after differentiating u= 3x ?

OpenStudy (anonymous):

i did e^3x = ln 3x

hartnn (hartnn):

e^3x doesn't equal ln 3x.... before starting integration, we must know basic derivatives ? if you forgot, please revise the derivatives formulas!

OpenStudy (anonymous):

u = 3x 1/3du=dx

OpenStudy (anonymous):

3x just equals 3 with derivatives

OpenStudy (anonymous):

but idk where to take this problem

hartnn (hartnn):

1/3 du = dx is correct :)

hartnn (hartnn):

so, we should change our integration variable from x to u

OpenStudy (anonymous):

ok so then e^u * 1/3 du

hartnn (hartnn):

yes! good :) now change the limits too! when x= 1, u=... ? when x=3, u=.... ?

OpenStudy (anonymous):

e^1 = 2.71 e^3 = 20.09

hartnn (hartnn):

........ 3x = u, so x is just u/3 ! right ? now try again :)

OpenStudy (anonymous):

ok 1/3 = .333 3/3 = 1

hartnn (hartnn):

correct, so now we have \(\large \int \limits_{1/3}^1e^udu/3=\large (1/3)\int \limits_{1/3}^1e^udu\) can you integrate e^u ??

OpenStudy (anonymous):

\[\int\limits_{1}^{3}e ^{3x}dx = e ^{u} = e ^{u}* 1/3 du \]

hartnn (hartnn):

constants can be taken out of integration :)

OpenStudy (anonymous):

= e^3x * 1/3 = e^3*1 *1/3 - e^3*3 * 1/3

OpenStudy (anonymous):

2694.33

OpenStudy (anonymous):

i did that backwards, meant to put the 3 - 1 not 1-3

hartnn (hartnn):

i don't understand what u did , there shouldn't be any x, once we changed the variable from x to u, there should just be 'u' \(\large \int \limits_{1/3}^1e^udu/3=\large (1/3)\int \limits_{1/3}^1e^udu=(1/3)[e^u]^1_{1/3}\) because integral of e^u du is just e^u

hartnn (hartnn):

\(\large (1/3) [e^1 -e^{1/3}]\) ^thats substituting the limits we can keep this as final answer or use calculator and find decimal approximation...

OpenStudy (anonymous):

wow this was hard....thank you

hartnn (hartnn):

with some practice, it'll become easy :) welcome ^_^